OpenStudy (anonymous):

evaluate integral: xe^(2x^2) dx

OpenStudy (anonymous):

use a u-sub of \(u = 2x^2\) and I think you'll get good results.

OpenStudy (anonymous):

du=4x^2.. what do I do from here?

OpenStudy (anonymous):

\[du \neq4x ^{2}\]

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

Finding the derivative is one of the simplest operations in calculus. What is the derivative of\[2x ^{2}\]

OpenStudy (anonymous):

yeah i know i have 4x

OpenStudy (anonymous):

\[du = 4x\ dx \]\[\implies dx = \frac{1}{4x}du\]\[\implies \int xe^{2x^2}dx \]\[= \int xe^udx = \int x e^u (\frac{1}{4x})du = \frac{1}{4}\int e^u du\]

OpenStudy (anonymous):

is that my final answer?